PA 4-4 (Algo) Mr. Ks is a very popular hair salon. It... Mr. Ks is a very popular hair salon. It offers high-quality hairstInstruction: Do not round your intermediate calculations. Round your final answers to 2 decimal places. a. What is the labor
Answer #1

a. labor content = 9+18+30 = 57 minutes

b.
Idle time at S1 => 30-9 = 21 minutes
Idle time at S2 = 30-18 = 12 minutes

total idle time = 21+12 = 33 minutes

average Labor utilization = (labor content/(labor content + total idle time)*100 = (57/(57+33))*100 = 63.33333333% = 63.33% (rounded to 2 decimal places)

c.

cost of direct labor = hourly wage/flow rate = (3*24)/(60/30) = 36 per customer

d.

Additional resource would help S3 shifting bottleneck to S2

cost of direct labor = hourly wage/flow rate = (3*24)/(60/18) = 21.6 per customer

e.

S3 will still be bottleneck however flow rate will increase

cost of direct labor = hourly wage/flow rate = (3*24)/(60/22) = 26.4 per customer

f.

one additional worker to be assigned to activities 4 and 5. Now, Additional resource would help S3 shifting bottleneck to S2.

process capacity = 60/18 = 3.333333333 customers per hour = 3.33 customers per hour (Rounded to 2 decimal places)

g.

It will still keep S3 as bottleneck with processing time per unit as 26 minutes

So process capacity = 60/26 = 2.307692308 = 2.31 customers per hour (Rounded to 2 decimal places)